find x and y
2x + 3y = 6
3x - 4y = -3
: find x and y
2x + 3y = 6
3x - 4y = -3

Answer by checkley77(3423)

2x + 3y = 6 multiply by 3
3x - 4y = -3 multiply by -2
6x+9y=18
-6x+8y=6 now add.
----------------------------
17y=24
y=24/17 answer.
2x+3*24/17=6
2x+72/17=6
2x=6-72/17
2x=(6*17-72)/17
2x=(102-72)/17
2x=30/17
x=(30/17)/2
x=30/17*1/2
x=30/34=15/17 answer.
proof:
3*15/17-4*24/17=6
45/17-96/17=-3
-51/17=-3
-3=-3